OpenStudy (mkapad01):

The power P (watts) of an electric circuit is related to the circuit's resistance R (ohms) and current I (amperes) by the equation P=RI^2. b) How is dR/dt related to dI/dt if P is constant?

OpenStudy (johnweldon1993):

Take your derivative \[\large \frac{dP}{dt} = (I^2\frac{dR}{dt}) + (2RI \frac{dI}{dt})\] If P is constant...we know dP/dt will = 0 so \[\large I^2\frac{dR}{dt} + 2RI\frac{dI}{dt} = 0\] Now just solve for \(\large \frac{dR}{dt}\)

OpenStudy (mkapad01):

awesome, thanks! :)